Meteorite Classifier Report from Liam Pedersen

Work here is progressing slowly, but steadily. There have been a number of technical mishaps that have delayed things. Nevertheless, they seem to be solved and I expect trials to begin within 3-4 days.

Pirrit Hills and Martin Hills Field Trip

We footsearched both these areas, finding one meteorite at Martin Hills. Inclement weather and the extreme cold made digital photagraphy of rocks and spectroscopy with the current equipment impossible. The digital camera would not operate, even though kept close to my body for warmth.

Instead, I concentrated on foot search and rock sample collection. Approximately 20 rock samples that appeared to cover all the rock types in the area (mostly granite) were collected and individually labelled. Color slide photographs were taken of some of these, along with photographs of the area and the boulder fields where we searched. These images should be useful to train the next classifier.

Spectra of the samples have been taken back in Patriot Hills base camp, and our currently being used with last seasons data set to retrain the classifier for the coming tests.

Hi-Resolution Pan/Tilt Mounted Color Camera

Code to compute the image scaling [mm/pixel] was NOT implemented in the camera-PTU code. This is essential for the classifier. To remedy this I have done experiments with Nomad, taking images of a calibration target at various distances from the vehicle and at various zoom encoder settings. These have been sufficient to model the camera system (panptilt unit, zoom lens + encoders, sony xc003 camera, and placement upon Nomad) with enough accuracy for our purposes. Progress was hampered by the disorganized and unfinished state of the code and lack of documentation. It has been necessary to rewrite large parts of the camera system code.

All the camera enclosures have suffered from interior condensation freezing on the front glass plate and obscuring the camera field of view. Cause appears to be small imperfections in the enclosures allowing snow to enter, and residual water vapour. I have duct taped the vandicam enclosure and inserted cloth bags with rice (best dessicant available here). That appears to have solved the problem.

Spectrometer

I have made some minor changes to the spectrometer system to cope with some problems with the attenuator that cropped up, possibly due to damage in transit. There is also a problem with an erratic connector, so the system is unstable in the off robot configuration (a different connector is used on the robot). I should like to compliment John Bates on producing good and well documented code for the spectrometer that made it much easier to diagnose and correct the problem.

To do

Prior to commencing trials need to:
  • Remount spectrometer on robot, and install code fix. Awaiting suitable weather to take off shell.
  • Retrain classifier (both image segmenter and bayes network) on latest data from region.
  • Add some other minor features to camera code that are still missing.
  • Baring further equipment breakdowns there appear to be no insoluble obstacles to testing in a few days.

    Expedition Report from William Cassidy


    Nov. 17, 1998, Tuesday: Patriot Hills Camp

    The wind blowing through camp was milder this morning and the lenticular clouds were still there. By evening the surroundings were almost windless, and while it seems colder than it has been recently, the weather is otherwise excellent. The lenticular clouds are gone from Independence Hills. Stewart spent the morning with Nicolas at Camp Crickett. Nicolas was taking images with his laser experiment and Stewart was maneuvering Nomad. Eventually his fingertips lost all feeling and he was forced to stop.

    Stewart, Christian, Fernando and Nicolas went to Nomad in the afternoon to take some more stereo camera photos with different filters. When Stewart opened the filter case, most of the filters blew away. Two polaroid filters were saved, so he took stereo photos in which one set of cameras has polaroid filters attached, and one does not. Eventually the computer in the robot gave problems, so he had to discontinue the operation. Nicolas carried out his own agenda of experiments.

    At Patriot Hills camp, Matt tried to download his data from Nomad via ARLAN, a wireless ethernet connection but lost it twice, forcing him to start again each time. It will require several hours to retreive the entire body of data, and he finally had to give up for the day. Mike and Liam mounted the spectrometer on Nomad, to be ready for Nomad's onward trip to the Independence Hills moraine. Sib went out to Nomad to turn it off and found it was out of fuel, so he refueled it, ready for the next day's operations.